### v2

# Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance

Abstract
While large language models (LLMs) have shown strong general reasoning capabilities, their effectiveness in financial reasoning—which is crucial for real-world financial applications—remains underexplored. In this study, we conduct a comprehensive evaluation of 24 state-of-the-art general and reasoning-focused LLMs across four complex financial reasoning tasks involving financial text, tabular data, and equations. We assess key capabilities such as numerical reasoning, tabular interpretation, financial terminology comprehension, long-context understanding, and equation-based problem solving. Our analysis reveals that while data quality and pretraining contribute to performance, general techniques like chain-of-thought (CoT) fine-tuning offer limited gains in financial tasks. To address this, we propose two domain-adapted models, Fino1-8B and Fino1-14B, trained with CoT fine-tuning and reinforcement learning using domain-specific reasoning paths. Our models are trained on a carefully curated dataset integrating high-quality examples from diverse sources, covering financial reports, tables, equations, and structured XBRL texts. Despite limited training data, they achieve an 7–9% performance improvement, outperforming several advanced LLMs—including GPT-o1, GPT-o3-mini, GPT-4.5, and comparable with DeepSeek models (V3 and R1)—demonstrating strong practical value in resource-constrained scenarios. Our findings highlight the need for domain-specific adaptations in financial reasoning, and we release all datasets, models, and code for future research [^1] https://github.com/The-FinAI/Fino1.
1 Introduction
Advancements in large language models (LLMs) have demonstrated remarkable performance across various natural language pocessing tasks, including content generationVan Veen et al. (2024); Zhang et al. (2024), language translation Xu et al. (2024), and sentiment analysis Xing (2024); Miah et al. (2024). More recently, reasoning-enhanced models such as OpenAI’s o1 and DeepSeek’s R1 have been developed to extend LLM capabilities in complex reasoning Guo et al. (2025); Jaech et al. (2024). These models exhibit significant improvements, particularly in mathematical and logical tasks that require complex reasoning Temsah et al. (2024); Zhong et al. (2024). However, despite their success in general reasoning, their performance in the financial domain remains largely unexplored.
Financial reasoning is essential in the financial domain, where professionals must extract key insights from lengthy reports and multi-table documents. It requires not only numerical computation but also a deep understanding of financial terminology, regulations, and economic concepts Xie et al. (2023, 2024a). Unlike general reasoning tasks, financial reasoning involves (1) domain-specific language comprehension, (2) reasoning over financial concepts and numbers, and (3) interpreting structured data such as financial tables Xie et al. (2024b). Evaluating LLMs in this context is critical for closing the gap between general AI reasoning and real-world financial applications.
To address this gap, we firstly comprehensively evaluate the performance of existing powerful reasoning models and conduct an in-depth analysis of their capabilities in financial tasks, to provide insights into the strengths and limitations of existing reasoning models in financial applications. We focus on the following key aspects: (1) Transferability of general-domain reasoning enhancements – Assessing whether reasoning-enhanced LLMs, originally optimized for general-domain reasoning, also demonstrate improved performance in financial reasoning tasks. (2) Impact across financial tasks – Analyzing how the reasoning enhancement process influences LLM performance across different types of financial tasks. (3) Scaling and performance gap – Investigating the differences between smaller reasoning models and large-scale LLMs in financial reasoning. (4) Limits of LLMs in complex financial tasks – Evaluating the extent to which LLMs can handle extremely complex financial reasoning challenges.
We evaluated 24 LLMs from the GPT Achiam et al. (2023), LLaMA Dubey et al. (2024), DeepSeek Liu et al. (2024), and Qwen Yang et al. (2024) families on four financial reasoning benchmarks: FinQA Chen et al. (2021), DM-Simplong Zhao et al. (2024a), XBRL-Math Wang et al. (2025) and DM-Complong. These datasets span financial texts, tables, and equations, enabling evaluation of models in numerical reasoning, domain-specific terminology comprehension, long-context processing, and equation-based reasoning tasks. Our evaluation encompasses both general-purpose and reasoning-enhanced models, ranging from smaller-scale models (e.g., LLaMA 3-8B-InstructDubey et al. (2024)) to high-capacity models (e.g., GPT-o1Jaech et al. (2024)), as well as reasoning-augmented models incorporating strategies such as data sampling and financial Chain-of-Thought (CoT) construction. Experimental results reveal three key insights: (1) Reasoning-enhanced models like DeepSeek-R1 and GPT-o1 do not consistently outperform general-purpose models on financial tasks, likely due to challenges related to domain-specific terminology and multi-table reasoning; (2) Performance gains from larger model sizes and improved pretraining tend to plateau beyond 32B or 70B parameters;
To develop reasoning models tailored for financial domain, we build novel financial reasoning-enhanced large language models Fino1-8B and Fino1-14B based on LLaMA-3.1-8B-InstructDubey et al. (2024) and Qwen2.5-14B-InstructYang et al. (2024), chosen for their strong performance among smaller-scale LLMs. To enhance reasoning capabilities, we apply CoT fine-tuning and reinforcement learning (RL), inspired by the training strategies of HuatuoGPT-o1 Chen et al. (2024). We leverage a diverse set of high quality financial reasoning datasets—including FinQA Chen et al. (2022), TATQA Zhu et al. (2021), DocMath (Simpshort and Compshort) Zhao et al. (2024b) , DocFinQA Reddy et al. (2024a), and BizBench-QA Krumdick et al. (2024)—covering tasks such as long-context question answering, complex reasoning over financial documents, and table-text integration. We construct high-quality reasoning paths using GPT-4o, combined with a novel data filtering process that selects more challenging samples for training—particularly for the 14B model—ensuring the model learns from difficult, domain-specific reasoning cases rather than generic QA pairs. Experimental results show that this targeted training strategy enables our models to surpass the performance of advanced closed-source models such as GPT-o3-mini and GPT-o1, and approach most powerful LLMs like DeepSeek-V3, DeepSeek-R1, and GPT-4o, demonstrating that domain-specific reasoning-enhanced training with limited, high-quality data can be highly effective.
To the best of our knowledge, this work presents the first comprehensive study of reasoning-enhanced LLMs for financial tasks involving diverse inputs such as financial documents, tables, equations, and structured XBRL texts. First, we conduct a systematic evaluation of 24 general-purpose and reasoning-enhanced LLMs across multiple financial reasoning benchmarks. Second, we introduce Fino1, the first suite of domain-specific, reasoning-enhanced LLMs, trained using GPT-4o-distilled reasoning paths with a novel data curation and filtering strategy. Third, our results show that domain-specific reasoning training with limited high-quality data consistently outperforms general-domain strategies and even surpasses larger closed-source models. We release all models, datasets, and code to support future research in financial AI.
2 Methods
2.1 Datasets
To evaluate the financial reasoning capabilities of reasoning-oriented and general LLMs, we select four diverse benchmark tasks. (1) FINQA Chen et al. (2021) is a large-scale dataset that emphasizes complex numerical reasoning in financial reports, requiring the integration of structured tables and unstructured text, along with a deep understanding of domain-specific terminology. (2) DocMath (Simplong, Complong) Zhao et al. (2024a) evaluates LLMs’ numerical reasoning abilities over long, specialized documents with multi-tiered tables, testing their capacity to extract and reason over extended contexts. The simplong subset contains relatively easier questions, while the complong subset includes more complex questions that require advanced reasoning across tables and context. (3) XBRL-Math Wang et al. (2025) targets reasoning over structured financial data in XBRL filings, which involve standardized taxonomies, hierarchical disclosures, and formulas, challenging models to interpret intricate numerical relationships.
Dataset Size Data Types Average Token
FinQA 1100 Tables and Texts 1,128
DM-Simplong 100 Tables and Texts 4,330
DM-Complong 300 Tables and Texts 39,983
XBRL-Math 90 Texts and Equations 397
Table 1: Overview of the datasets used in the study.
More details of the datasets are listed in Table 1. By conducting experiments on these three datasets, we aim to evaluate the model’s ability in various aspects, including understanding financial terms, extracting relevant numbers and entities from financial reports from different sources, and reasoning over long contexts and multiple tables.
2.2 Evaluated Models
[…]
The evaluated models are summarized as follows: GPT-4o Hurst et al. (2024): OpenAI’s flagship model with multimodal capabilities and improved real-time reasoning. GPT-o1 Jaech et al. (2024): A reasoning-optimized variant in the GPT series with strong step-by-step problem-solving ability. GPT-o3-mini: A compact version of GPT-o3, optimized for efficiency with robust language understanding. DeepSeek-V3 Liu et al. (2024): A 671B Mixture-of-Experts (MoE) model with 37B active parameters per token, designed for efficient reasoning. DeepSeek-R1 Guo et al. (2025): A first-generation reasoning model with competitive performance on math, code, and reasoning tasks. Qwen2.5-72B-Instruct Yang et al. (2024): A strong multilingual model with advanced instruction tuning. Qwen2.5-72B-Instruct-Math: A math-specialized variant fine-tuned for complex problem solving. LLaMA-series Dubey et al. (2024): Widely adopted open-source models (8B–70B) with strong instruction-tuned variants used in our experiments. DeepSeek-R1-Distill Guo et al. (2025): Distilled models from DeepSeek-R1, fine-tuned using LLaMA and Qwen backbones to balance performance and efficiency. Other reasoning-enhanced models: Includes S1-32BMuennighoff et al. (2025), Limo Ye et al. (2025), and FinR1-7B Liu et al. (2025b), incorporating reasoning strategies such as financial CoT construction and task-specific fine-tuning.
Model Name Parameters Enhanced for Reasoning Tasks Context Window Size Close/Open Source Reasoning Enhanced Training Data
GPT-4o Unknown No 128k Closed -
GPT-o1 Unknown Yes 128k Closed Public and proprietary data (Human-annotated CoT, MCTS-assisted Synthetic data)
GPT-o3-mini Unknown Yes 128k Closed Public and proprietary data
DeepSeek-V3 671B No 128k Open -
DeepSeek-R1 671B Yes 128k Open Cold-start data generation, post-processing data
GPT-4.5 Unknown No 128k Closed -
Qwen2.5-72B-Instruct 72B No 128k Open -
Qwen2.5-72B-Instruct-Math 72B Yes 128k Open Synthetic data from Qwen, high-quality mathematical data, CoT, TIR
DeepSeek-R1-Distill-Llama-70B 70B Yes 128k Open Distilled from R1
Llama3-70B-Instruct 70B No 8k Open -
Llama3.1-70B-Instruct 70B No 128k Open -
Llama3.3-70B-Instruct 70B No 128k Open -
Qwen2.5-32B-Instruct 32B No 128k Open -
DeepSeek-R1-Distill-Qwen-32B 32B Yes 128k Open Distilled from R1
Limo 32B Yes 128k Open Public and proprietary data with Qwen filteration
S1-32B 32B Yes 128k Open Public and proprietary data + original data, filtered by quality, difficulty, diversity
Qwen/QwQ-32B 32B Yes 128k Open -
Qwen2.5-14B-Instruct 14B No 128k Open -
DeepSeek-R1-Distill-Qwen-14B 14B Yes 128k Open Distilled from R1
DeepSeek-R1-Distill-Llama-8B 8B Yes 128k Open Distilled from R1
Llama3-8B-Instruct 8B No 8k Open -
Llama3.1-8B-Instruct 8B No 128k Open -
Qwen2.5-7B-Instruct 7B No 128k Open -
FinR1-7B 7B Yes 128k Open 60k CoT data for financial tasks
Table 2: Summary of evaluated large language models, including their parameter sizes, reasoning capabilities, input limits, source availability, and reasoning enhancement strategies. MCTS refers to Monte Carlo Tree Search. CoT refers to chain-of-thought reasoning. TIR means tool-integrated reasoning.
2.3 Evaluation Settings
In each dataset, LLMs are presented with a context containing financial reports, tables, equations, and symbols, along with a question. Following prior work Zhao et al. (2024a), we use a generic prompt (see appendix D for more details) to instruct models to answer based on the given context. Models must extract relevant information and perform mathematical reasoning using numbers and concepts from the context. While outputs are typically in mathematical form, variations in formats (e.g., percentages, rounding) may cause differences. Reasoning-focused models like DeepSeek-R1 and GPT-o1 often generate step-by-step reasoning. Inspired by DocMath Zhao et al. (2024a), we adopt an LLM-as-judge evaluation approach: GPT-3.5-Turbo extracts final answers, followed by a mathematical comparison to assess correctness.
For large models (e.g., DeepSeek-V3, DeepSeek-R1, GPT series), we run evaluations via APIs from TogetherAI [^2] https://www.together.ai and OpenAI [^3] https://platform.openai.com, using versions GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06. For models up to 72B parameters, we use the VLLM framework on 8 A100 GPUs (80GB each), with a max generation length of 512 tokens.
2.3.1 Reasoning Data Curation
To enhance domain-specific reasoning capabilities in LLMs, we curate a high-quality training dataset by integrating diverse financial reasoning benchmarks. Our dataset combines a wide range of reasoning tasks—covering numerical computation, table-text integration, cross-table correlation, and long-context understanding—reflecting the complexity of real-world financial documents.
Specifically, we incorporate data from several well-established benchmarks, including FinQA Chen et al. (2022), DocFinQA Reddy et al. (2024b), TAT-QA Zhu et al. (2021), DocMath-Eval Zhao et al. (2024b), and BizBench-QA Krumdick et al. (2024). These datasets collectively provide rich and diverse coverage across task types and difficulty levels, supporting the development of robust financial reasoning models. FinQA provides QA pairs combining textual and tabular financial data; DocFinQA extends this to long-context documents; TAT-QA focuses on complex numerical reasoning; DocMath-Eval (CompLong and SimpShort subsets) targets reasoning over multi-table and single-table scenarios; and BizBench-QA enriches the corpus with code-based financial reasoning tasks. Please see Appendix B for detailed descriptions of each dataset.
2.4 Data Collection and Reasoning Path Construction
To ensure a balanced dataset of both short and long context samples, we collect all data from the FinQA train set, TATQA, DocMath-CompLong, and DocMath-SimpShort. Additionally, we sample 1,000 examples each from BizBench-QA and DocFinQA to balance short and long reasoning tasks. In total, we gather 9,851 financial QA samples for reasoning path construction. Inspired by Ye et al. (2025); Muennighoff et al. (2025), which suggest that simple QA pairs may not effectively improve reasoning capabilities, we apply a filtering strategy to retain more challenging examples. Specifically, we use Qwen-14B-Instruct to generate answers for each question and employ GPT-3.5-Turbo as a verifier. QA pairs correctly answered by Qwen-14B are removed, resulting in a more difficult training subset of 3,472 QA pairs for optimizing more advanced models.
2.5 Reasoning Path Generation and Refinement
For reasoning path generation, we use GPT-4o, chosen for its strong performance across datasets (Table 3) and its stability in generating high-quality outputs. Following the CoT refinement framework from HuatuoGPT-o1 Chen et al. (2024), we employ a verifier-based iterative process to refine the reasoning paths generated by GPT-4o. If a reasoning path is rejected by the verifier, refinement is guided by one of the following strategies: (1) Backtracking: Revisiting previous steps to correct logical inconsistencies. (2) Exploring New Paths: Generating alternative reasoning routes when the current path fails. (3) Verification: Cross-checking intermediate steps to ensure logical soundness. (4) Correction: Making targeted adjustments to rectify specific errors. The CoT refinement process begins with an initial prompt to generate a reasoning path and final answer. If the verifier identifies the answer as incorrect, one of the above strategies is applied iteratively until a correct and logically consistent answer is produced.
2.5.1 Model Training
We develop our financial reasoning models based on curated reasoning paths, focusing on smaller-sized models for efficient real-world deployment. We select LLaMA-3.1-8B-Instruct and Qwen2.5-14B-Instruct as backbone models, given their strong performance among small and mid-sized LLMs. We adopt a two-stage training approach. In the first stage, we perform Supervised Fine-Tuning (SFT), using questions as input and reasoning paths with final answers as output. For the 8B model, we use only FinQA reasoning paths to assess the impact of domain-specific fine-tuning. For the 14B model, we use the full set of sampled and filtered data to leverage its greater capacity for diverse and complex reasoning. In the second stage, we apply RL using a verifier-based sparse reward mechanism and optimize with Proximal Policy Optimization (PPO), following Chen et al. (2024). A reward model evaluates the correctness and coherence of the reasoning path. If errors are found, the model iteratively refines its outputs through backtracking, alternative path exploration, verification, and correction, improving logical consistency and equation-solving accuracy. During SFT, models are trained for three epochs with a learning rate of 5e-6 and a maximum sequence length of 8192 tokens. LoRA is applied for efficient parameter updates. In the RL stage, training continues for three epochs with a reduced learning rate of 5e-7, guided by the reward model to refine reasoning quality via PPO.
3 Results
Table 3 presents the performance of various LLMs across four financial reasoning benchmarks: FinQA, DM-Simplong, XBRL-Math, and DM-Complong. Our models, Fino1-8B and Fino1-14B, demonstrate strong and consistent results across tasks, outperforming several larger and closed-source models. Fino1-14B achieves an average score of 60.25, surpassing many larger-scale models such as LLaMA-3.3-70B-Instruct (56.04), Qwen2.5-72B-Instruct (53.71), GPT-o1-preview (54.05), GPT-o3-mini (57.89), and even GPT-4.5 (60.43). It also achieves the best performance among all models on two benchmarks—DM-Simplong (60.00) and XBRL-Math (86.67). These results demonstrate that high-quality, domain-specific reasoning training can enable smaller models to rival—and even surpass—the performance of significantly larger and proprietary LLMs.
Fino1-8B also performs competitively with an average score of 50.77, surpassing other 8B models and even some 14B and 32B models. Both models excel on XBRL-Math, where Fino1-14B achieves the top score of 86.67. Moreover, Fino1-14B and Fino1-8B significantly outperform FinR1-7B, a financial reasoning-enhanced model trained on a much larger 60K task-aligned dataset, demonstrating that larger data volume does not necessarily lead to better performance, and that carefully curated, high-quality reasoning data—even in smaller quantities—can yield superior results.
Models FinQA DM-Simplong XBRL-Math DM-Complong Average
GPT-4o 72.49 60.00 72.22 39.33 61.01
GPT-o1-preview 49.07 56.00 74.44 36.67 54.05
GPT-o3-mini 60.87 59.00 76.67 35.00 57.89
DeepSeek-V3 73.20 53.00 76.67 42.33 61.30
DeepSeek-R1 65.13 53.00 86.67 38.67 60.87
GPT-4.5 68.94 59.00 74.44 39.33 60.43
Meta-Llama-3-70B-Instruct 58.92 41.00 56.67 13.67 42.57
Llama-3.1-70B-Instruct 63.18 48.00 63.33 34.33 52.21
Llama-3.3-70B-Instruct 68.15 54.00 70.00 32.00 56.04
Qwen2.5-72B-Instruct 73.38 59.00 67.78 14.67 53.71
Qwen2.5-Math-72B-Instruct 69.74 42.00 83.33 5.00 50.02
DeepSeek-R1-Distill-Llama-70B 66.73 53.00 86.67 30.67 59.27
Qwen2.5-32B-Instruct 73.11 56.00 65.56 30.00 56.17
Qwen/QwQ-32B 61.22 46.00 84.44 20.00 52.92
DeepSeek-R1-Distill-Qwen-32B 65.48 55.00 84.44 24.67 57.40
Limo 63.44 45.00 61.11 15.33 46.22
S1-32B 66.81 53.00 84.44 24.00 57.06
Qwen2.5-14B-Instruct 67.44 59.00 57.78 26.67 52.72
DeepSeek-R1-Distill-Qwen-14B 63.27 44.00 84.44 21.00 53.18
DeepSeek-R1-Distill-Llama-8B 45.96 33.00 81.11 15.67 43.94
Meta-Llama-3-8B-Instruct 41.97 29.00 48.89 6.00 31.47
Llama-3.1-8B-Instruct 54.13 34.00 62.22 14.30 41.16
Qwen2.5-7B-Instruct 55.37 41.00 42.22 17.67 39.07
FinR1-7B 58.74 37.00 30.00 13.67 34.85
Fino1-8B 60.87 40.00 82.22 20.00 50.77
Fino1-14B 70.01 60.00 86.67 24.33 60.25
Table 3: Performance of different LLMs on four financial reasoning datasets.
Beyond overall performance, the comparison across models provides additional insights that can guide the development of LLMs for financial applications:
(1) General reasoning enhancements do not necessarily improve financial task performance. Models like GPT-o1 and DeepSeek-R1, though strong in general reasoning, underperform on financial benchmarks, highlighting the importance of domain-specific knowledge and structured numerical reasoning. Moreover, reasoning enhancement strategies based on general-domain data, as seen in S1-32B and Limo, often fail to improve and sometimes even degrade performance compared to their backbone models (e.g., Qwen2.5-32B). These results underscore that high-quality, domain-aligned reasoning paths are more effective than general reasoning augmentation, and tailored financial training is crucial for optimal model performance.
(2) Reasoning enhancements vary in effectiveness across financial tasks. Models like GPT-o1 perform worse than GPT-4o on FinQA (49.07 vs. 72.49) and DM-Simplong (56.00 vs. 60.00). Qwen2.5-Math improves in XBRL-Math (67.78 $\rightarrow$ 83.33) but drops sharply in FinQA (69.74 $\rightarrow$ 42.00) and DM-Simplong (59.00 $\rightarrow$ 42.00). DeepSeek-R1-Distill-Llama-8B also excels in XBRL-Math (81.11) but shows limited gains elsewhere. These results suggest that strategies like CoT or math fine-tuning may offer benefits for structured tasks but often fail to generalize effectively to broader financial reasoning. Compared with Qwen2.5-72B-Instruct and its math enhanced version, we can see that while TIR and synthetic building appears to help in equation-heavy tasks like XBRL-Math, its impact on other types of financial tasks is less clear and may be influenced by multiple factors such as task structure, data modality, or domain complexity.
(3) Larger model size offers limited benefits in financial reasoning. While larger models typically excel in general NLP tasks, their gains in financial tasks are not proportional. For example, DeepSeek-70B (66.73 FinQA, 53 DM-Simplong, 86.67 XBRL-Math) performs similarly to the smaller DeepSeek-R1-Distill-Qwen-32B (65.48, 55, 84.44), despite having less than half the parameters. Across models like Qwen2.5-72B, DeepSeek-R1-Distill-Llama-70B, and GPT-4o, performance tends to plateau around 70B. This may reflect the structured nature of tasks like XBRL-Math, where arithmetic precision outweighs complex language reasoning, enabling smaller models to compete effectively.
(4) Improved pre-training data and strategies in general domains also benefit financial tasks. Llama 3.1-8B-Instruct shows a notable performance boost over Llama 3-8B-Instruct (avg. 50.12 vs. 39.95), with gains in XBRL-Math (62.22 vs. 48.89) and FinQA (54.13 vs. 41.97). Similarly, Llama 3.1-70B-Instruct outperforms Llama 3-70B (58.17 vs. 52.20), especially in DM-Simplong (48.00 vs. 41.00), while Llama 3.3-70B-Instruct achieves the highest overall performance (64.05), driven by improvements in FinQA (68.15) and XBRL-Math (70.00). These gains stem from larger, more diverse training corpora (e.g., 15T+ tokens in Llama 3.1) and enhanced post-training techniques like online preference optimization. Such strategies improve general reasoning while also boosting performance in domain-specific tasks like financial reasoning.
(5) Existing reasoning models struggle with long contexts and cross-table correlations. Financial tasks often require integrating information from lengthy reports, structured tables, and equations, yet current models show clear limitations. For example, GPT-4o performs well on FinQA (72.49) but drops on DM-Simplong (60.00), indicating challenges in combining textual and symbolic reasoning. Models like DeepSeek-R1 and DeepSeek-R1-Distill-Qwen-32B excel in XBRL-Math but fail to generalize. Performance on DM-Complong remains low across models (e.g., LLaMA-3.3-70B: 32.00). Smaller models like DeepSeek-R1-Distill-Llama-8B also perform poorly with 33.00 for DM-Simplong. These results highlight the need for improved long-context modeling and structured data integration in financial reasoning.
(6) Financial reasoning enhancement improves performance across all tasks. Compared to LLaMA 3.1-8B-Instruct (41.16), DeepSeek-R1-Distill-Llama-8B shows a modest gain (43.94). In contrast, our Fino1-8B, trained on GPT-4o-distilled reasoning paths with backtracking strategies, achieves a substantial improvement, reaching an average score of 50.77, and consistently outperforms baselines on FinQA (60.87 vs. 54.13), DM-Simplong (40.00 vs. 34.00), and XBRL-Math (82.22 vs. 62.22). Fino1-14B further boosts performance, achieving an average score of 60.25, outperforming other 14B models (e.g., Qwen2.5-14B: 52.72) and even rivaling state-of-the-art models like DeepSeek-R1 (60.87) and GPT-4o (61.01). Notably, although Fino1-8B and Fino1-14B were trained with limited financial CoTs, both models show improvements across all tasks, highlighting the effectiveness of high-quality reasoning paths and structured training strategies such as backtracking, path exploration, and verification.
Overall, these insights highlight that financial reasoning poses unique challenges distinct from general NLP and mathematical reasoning. The performance variability across tasks underscores the need for specialized training strategies and architectural improvements to better support financial comprehension, structured data processing, and multi-step numerical reasoning.
4 Conclusion
This study presents a comprehensive evaluation of 24 general-purpose and reasoning-augmented LLMs on financial reasoning tasks, encompassing financial text understanding, numerical computation, and structured data interpretation. Our findings show that existing models, despite strong performance on general reasoning benchmarks, often fail to generalize to financial domains—particularly in handling long contexts, multi-table reasoning, and domain-specific terminology. To bridge this gap, we curated a high-quality, diverse financial reasoning dataset and introduced Fino1, a suite of domain-adapted models trained with GPT-4o-generated reasoning paths using chain-of-thought fine-tuning and reinforcement learning. Despite being trained on limited data, Fino1 models outperform several larger and advanced LLMs, including GPT-o1, GPT-o3-mini, GPT-4.5, and DeepSeek models (V3 and R1). These results highlight the effectiveness of domain-specific reasoning strategies and demonstrate the practical value of tailored model adaptation for real-world financial applications.
Appendix A Related Work
Our work builds upon three main research directions: the development of general-purpose and reasoning-enhanced language models, applications of LLMs in finance, and studies on reasoning capabilities in financial tasks.
A.1 General-Purpose and Reasoning-Enhanced LLMs
Recent years have witnessed significant advancements in large language models which demonstrated remarkable capabilities across various tasks, including natural language understanding, generation, and complex reasoning. Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting Wei et al. (2022), self-reflection Zhang et al. (2023), and reinforcement learning from human feedback Ouyang et al. (2022). These models have shown superior performance in mathematical reasoning Lewkowycz et al. (2022), symbolic manipulation Drori et al. (2023), and logical inference tasks Chen et al. (2023).
A.2 LLMs in Financial Applications
The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis Yang et al. (2020), market sentiment prediction Yang et al. (2023), and automated trading strategies Kou et al. (2024). Recent work has focused on developing finance-specific models like BloombergGPT Wu et al. (2023), FinGPT Yang et al. (2023) and PIXIU Xie et al. (2023), OpenFinLLMs Xie et al. (2024b), which are pretrained on financial corpora to better understand domain-specific terminology and concepts. These models have demonstrated effectiveness in tasks such as financial sentiment analysis, earnings call analysis, and regulatory compliance checking.
A.3 Reasoning Capabilities in Financial Tasks
Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference. Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents Chen et al. (2021), mathematical problem-solving in financial contexts Srivastava et al. (2024), and structured data interpretation Liu et al. (2025a). These works highlight the importance of both domain expertise and reasoning abilities in financial applications. Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities.
Appendix B Details Data Source
We provide here detailed descriptions of the datasets used to construct our training and evaluation corpus for financial reasoning enhancement.
  • FinQA Chen et al. (2022): FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables. It integrates both textual and tabular data, covering numerical reasoning tasks such as arithmetic operations, comparison, and multi-step calculations, making it well-suited for training reasoning-enhanced models.
  • DocFinQA Reddy et al. (2024b): An extension of FinQA designed for long-context reasoning. It pairs FinQA-style questions with full-length financial filings, increasing the average context length from under 700 words to approximately 123,000 words. This dataset better simulates real-world document scale and complexity, enabling robust evaluation of retrieval and reasoning over long financial documents.
  • TAT-QA Zhu et al. (2021): A large-scale financial QA dataset created from real financial reports. It combines both textual and tabular data and emphasizes diverse numerical reasoning tasks, including arithmetic operations, counting, sorting, and comparison. TAT-QA reflects real-world challenges in financial document interpretation.
  • DocMath-Eval Zhao et al. (2024b): This benchmark suite evaluates mathematical reasoning in domain-specific contexts. We utilize two subsets:
    • DocMath-CompLong: Constructed from quarterly and annual reports, it involves reasoning across long documents and multiple tables, presenting a challenging multi-step numerical reasoning setting.
    • DocMath-SimpShort: Built from FinQA and TAT-QA, this subset focuses on localized numerical reasoning tasks over short documents containing a single table.
  • BizBench-QA Krumdick et al. (2024): Contains 14,400 financial QA pairs across multiple task types, including SEC-NUM, CodeFinQA, CodeTAT-QA, and FinCode. These tasks span code generation, numerical span identification, and financial mathematical reasoning from various perspectives, enriching the training corpus with broader financial task coverage.
Figure 1: Error case 1.
Figure 2: Error case 2.
Appendix C Error Analysis
To further investigate why reasoning-enhanced models like DeepSeek-R1 and GPT-o1 underperform in financial tasks compared to general-purpose models such as GPT-4o and DeepSeek-V3, we conduct an error analysis using the FinQA dataset since these reasoning-enhanced models exhibit a significant performance drop on FinQA.
We focus on cases where DeepSeek-R1 produces incorrect answers while DeepSeek-V3 generates correct ones. Two common error patterns emerge. First, DeepSeek-R1 tends to over-reason and does not strictly adhere to instructions, as illustrated in Figure 1. In this case, the question requires a sum, but the model provides excessive details by breaking down the calculation instead of directly summing the values as instructed.
Second, compared to DeepSeek-V3, DeepSeek-R1 appears to lack financial sensitivity. For instance, in Figure 2, it fails to recognize critical financial nuances, leading to an incorrect response. Specifically, DeepSeek-R1 misinterprets the concept of "average share price" in the UK Employee Share Purchase Plan by incorrectly identifying it as the employee purchase price (\$4,078 per share) rather than the actual compensation expense per share (\$719 per share). This mistake stems from an inability to correctly differentiate between compensation expense per share and the full fair market value (FMV) calculation, resulting in an overestimated figure. In contrast, DeepSeek-V3 correctly understands that the compensation expense per share represents the direct financial impact per issued share, allowing it to provide the accurate answer. This highlights a fundamental gap in DeepSeek-R1’s financial reasoning capabilities, particularly in recognizing the accounting conventions used to report share-based compensation.
Limitations
Our study has several limitations that highlight areas for future improvement. First, our model scale is limited, as we only fine-tuned an 8B model (Fino1), while larger models (e.g., 70B) could potentially benefit more from reasoning enhancements. Second, our evaluation scope is restricted, covering only three financial reasoning tasks (FinQA, DM-Simplong, and XBRL-Math), which do not fully capture the breadth of financial NLP applications such as forecasting, financial sentiment analysis, and fraud detection. Third, our fine-tuning relies on a single dataset, FinQA, for reasoning path construction, limiting the model’s exposure to different financial reasoning patterns; incorporating additional datasets could improve generalization. Fourth, our reasoning path construction approach is simplified, as we generate paths using a single method (GPT-4o), whereas exploring multiple reasoning path generation strategies—such as ensemble approaches or human-annotated paths—could lead to more robust financial reasoning capabilities.
C.1 Future Direction
Developing effective reasoning models for financial tasks poses distinct challenges compared to general-domain reasoning. Our findings suggest three key directions for improvement: (1) Enhancing financial knowledge and terminology understanding through domain-specific pretraining using financial corpora, structured reports, and regulatory documents; (2) Improving multi-table reasoning and long-context comprehension by developing strategies that better infer logic across tables and optimize memory mechanisms for processing lengthy financial documents; and (3) Refining reasoning-enhanced strategies, as current approaches like GPT-o1’s chain-of-thought prompting and DeepSeek-R1’s self-reflection via reinforcement learning yield mixed results—highlighting the need for more tailored reasoning strategies to address both structured numerical and unstructured textual financial reasoning.
Ethics Statement
While our study demonstrates significant advancements in financial reasoning with LLMs, Fino1 remains a research prototype and is not yet suitable for real-world financial applications. Our model inherits the well-documented limitations of large language models, including hallucinations, sensitivity to input phrasing, and potential biases in financial data sources, which could affect reliability in critical financial contexts. Furthermore, Fino1 has not undergone rigorous testing for high-stakes financial decision-making, such as investment strategies, regulatory compliance, or risk assessment, where incorrect outputs could lead to severe consequences. Additionally, the model’s reliance on a single dataset (FinQA) and a single reasoning path construction method (GPT-4o) limits its robustness and adaptability to the complex and evolving nature of financial environments. As with all AI-driven financial models, careful human oversight remains essential, and future work must focus on enhancing factual consistency, mitigating biases, improving domain adaptation, and ensuring alignment with regulatory frameworks before such models can be considered for real-world financial deployment.
Appendix D Prompt
Prompt Template for DocMath-CompLong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context. You need to first think through the problem step by step, documenting each necessary step. Then you are required to conclude your response with the final answer in your last sentence as Therefore, the answer is {final answer}’. The final answer should be a numeric value. ###Context <Full document here> ### Input <the question based on the document> Lets think step by step to answer the given question. ### Output
Prompt Template for DocMath-Simplong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context. You need to first think through the problem step by step, documenting each necessary step. Then you are required to conclude your response with the final answer in your last sentence as Therefore, the answer is {final answer}’. The final answer should be a numeric value. ###Context <simple long context here> ### Input <the question based on the context> Lets think step by step to answer the given question. ### Output
Prompt Template for FinQA task Please answer the given financial question based on the context. Context: <context> Question: <the question based on the context> Answer:
Prompt Template for XBRL-Math task You are a financial expert tasked with carefully reading, analyzing, and answering the following eXtensible Business Reporting Language. Please follow the steps below: INPUT: Read the eXtensible Business Reporting Language (XBRL) question: <XBRL related question>, formula: <the formula related to this question>, and the explanation: <term explanation>. Provide only the final answer which is the numerical result of the calculation. For formulas like ROI, provide percentages. Never use the percent symbol in percentages. OUTPUT: