### v3

# Fin-o1: On the Transferability of Reasoning-Enhanced LLMs and Reinforcement Learning to Finance

Abstract
As the fundamental capability behind decision-making in finance, financial reasoning poses distinct challenges for LLMs. Although reinforcement learning (RL) have boosted generic reasoning, the progress in finance is hindered by the absence of empirical study of building effective financial chain-of-thought (CoT) corpus, a systematic comparison of different RL methods, and comprehensive benchmarks. To address these gaps, we introduce FinCoT, the first open high-fidelity CoT corpus for finance, distilled from seven QA datasets by a novel three-stage pipeline that incorporates domain supervision, iterative LLM refinement, and difficulty-aware filtering. Based on FinCoT, we develop Fin-o1, the first open financial reasoning models trained via supervised fine-tuning and GRPO-based RL. Our models outperform existing financial reasoning models and SOTA general models such as GPT-o1, DeepSeek-R1, and GPT-4.5. We also investigate the effectiveness of three different RL methods in improving domain-specific reasoning, offering the first such empirical study. We finally propose FinReason, the first financial reasoning benchmark covering multi-table analysis, long-context reasoning, and equation-based tasks, and evaluate 29 LLMs. Our extensive experiments reveal general reasoning models excel on standard benchmarks yet exhibit obvious performance degradation in financial contexts; even finance-tuned models like Dianjin-R1 and FinR1 degrade on lengthy documents. In contrast, our Fin-o1 models consistently outperform their backbones and larger GPT-o1 and DeepSeek-R1, confirming the effectiveness of our data building and model training strategy. Our study further shows that GRPO yields reliable gains whereas PPO and DPO do not, highlighting the need for targeted data and optimisation rather than scale alone. These findings show that financial reasoning has unique challenges unmet by mere model scaling or generic techniques, highlighting the need for specialized data, models, and optimization. We release all datasets, models, and codes to support future research. 111 Dataset: https://huggingface.co/datasets/TheFinAI/FinCoT
      Fin-o1: https://huggingface.co/TheFinAI/Fin-o1-8B
                https://huggingface.co/TheFinAI/Fin-o1-14B
      Code: https://github.com/The-FinAI/Fino1
1 Introduction
Financial reasoning is a critical yet underexplored capability for large language models (LLMs), especially as they are increasingly deployed in high-stakes financial applications. Unlike general reasoning tasks, such as mathematical problem solving or code generation, financial reasoning requires models to interpret domain-specific language, analyze structured data like financial tables xie2024open , and extract insights from lengthy, complex documents reddy2024docfinqa . It demands not only numerical computation, but also contextual understanding of financial regulations, economic concepts, and ambiguous textual information xie2023pixiu ; xie2024finben .
Recent advances in reinforcement learning (RL) methods, including proximal policy optimization (PPO) schulman2017proximal , direct preference optimization (DPO) rafailov2023direct , and generalized reinforcement preference optimization (GRPO) tang2024generalized , have substantially enhanced the reasoning capabilities of large language models. These RL-based approaches form the foundation of state-of-the-art (SOTA) reasoning models such as OpenAI’s o1 jaech2024openai and DeepSeek’s R1 guo2025deepseek , achieving notable improvements on mathematical and logical problem-solving benchmarks temsah2024openai ; zhong2024evaluation . By directly optimizing for reasoning performance, RL strategies have become essential for advancing the complex reasoning abilities of LLMs.
However, improving LLMs with RL for financial reasoning is yet hindered by three critical challenges (As shown in Table 1). First, the absence of high-quality, publicly available chain-of-thought (CoT) datasets for financial reasoning limits the ability to model domain-specific reasoning processes. Existing resources are either proprietary jaech2024openai or adapted from general domains guo2025deepseek ; yang2024qwen2 , failing to capture the nuances of financial analysis. No empirical study has asked which source documents or path-generation strategies actually produce effective financial CoT traces. Second, while RL methods have enhanced general reasoning, their effectiveness in financial tasks remains unexplored, with no open-source models demonstrating domain-adapted improvements. The impact of different RL algorithms on financial reasoning remains untested: RL methods like PPO, DPO and GRPO have never been compared side-by-side, leaving practitioners without guidance on the best optimiser for the domain. Third, there is no comprehensive benchmark to systematically evaluate financial reasoning across key challenges such as domain-specific language understanding xie2024open , multi-table reasoning zhao2024docmath , long-context document analysis reddy2024docfinqa , and equation-based reasoning ma2023xbrl . Existing benchmarks like PIXIU xie2023pixiu , FinBen xie2024finben , and BizBench krumdick2024bizbench address only parts of this space. Moreover, current studies only evaluated a few LLM’s financial reasoning ability.
Table 1: Comparison of Existing Reasoning LLMs and Financial Benchmarks. SFT indicates super- vised fine-tuning. RL specifies the reinforcement learning method used. CoT data is categorized into Guide, indicate the presence of gold guidelines to direct reasoning generation, IterR means iteratively verification and refinement, DiffF indicates if involving difficulty-aware filtering to emphasize complex cases. Evaluation indicates whether the benchmarks involve single-table or multi-table reasoning, long-context analysis, and equation-based reasoning. ModN indicates the number of models evaluatd in each work. *Indicates works released after Fin-o1.
Model / Benchmark Domain CoT Data Model Evaluation
Public Guide IterR CompF SFT RL Single Multi Long Context Equations ModN
PIXIU xie2023pixiu Financial $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ 9
FinBen xie2024finben Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ 21
BizBench krumdick2024bizbench Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ 16
DocFinQA reddy2024docfinqa Financial $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ 12
S1-32B muennighoff2025s1 General $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 12
Limo ye2025limo General $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ 6
HuatuoGPT-o1 chen2024huatuogpt Other $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ PPO $\times$ $\times$ $\times$ $\times$ 10
DeepSeek-R1 guo2025deepseek General $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ GRPO $\times$ $\times$ $\times$ $\times$ $\times$
OpenAI-o1 jaech2024openai General $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Fin-R1finr1 * Financial $\times$ $\times$ $\times$ $\times$ $\checkmark$ GRPO $\checkmark$ $\times$ $\times$ $\times$ 10
DianJin-R1zhu2025dianjin * Financial $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ GRPO $\checkmark$ $\times$ $\times$ $\times$ 12
Fin-o1 (Ours) Financial $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ GRPO, PPO, DPO $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 29
To address these gaps, we present three key contributions. First, we curate FinCoT, the first high-quality, publicly available financial CoT dataset built by a novel three-stage framework that combines domain-aware supervision, iterative LLM refinement, and difficulty-aware filtering. We construct FinCoT from diverse sources including FinQA chen2022finqadatasetnumericalreasoning , ConvFinQA chen2022convfinqa , DocFinQA reddy2024docfinqa , TATQA zhu2021tatqaquestionansweringbenchmark , Econ_Logic quan2024econlogicqa , DocMath-Eval zhao2024docmath , and BizBench-QA krumdick2024bizbench . Our new framework yields high-fidelity financial CoT data that improves 8B and 14B models beyond SOTA reasoning models such as GPT-o1, DeepSeek-R1, and GPT-4.5. Second, we develop Fin-o1 (including Fin-o1-8B and Fin-o1-14) based on FinCoT, the first open-source financial reasoning large language models, trained via a two-phase framework combining supervised fine-tuning and RL-based reasoning enhancement. Fin-o1 consistently outperform existing financial reasoning models and SOTA general-reasoning LLMs such as GPT-o1, DeepSeek-R1, GPT-4.5. We systematically investigate the effectiveness of PPO schulman2017proximal , DPO rafailov2023direct , and GRPO tang2024generalized in improving domain-specific reasoning, offering the first empirical study on RL methods for financial reasoning. Third, we introduce FinReason, the first comprehensive benchmark designed to systematically evaluate the financial reasoning capabilities of LLMs. FinReason covers most LLMs with diverse tasks involving domain-specific language, structured data, and long-context documents, providing a robust framework on datasets such as FinQA chen2021finqa , DM-Simplong zhao2024docmath , XBRL-Math wang2025finnlp , and DM-Complong zhao2024docmath . We conduct a comprehensive evaluation on FinReason, covering 29 large language models from general-purpose, reasoning-enhanced, and financial-specific families, including GPT, LLaMA, DeepSeek, Qwen, and our Fin-o1 models.
Our results reveal that general reasoning models, despite their strong performance on generic benchmarks, struggle with financial reasoning tasks that require precise domain understanding, structured data interpretation, and long-context document reasoning. Frontier financial reasoning LLMs such as Dianjin-R1 and FinR1 exhibit performance degradation compared to their backbones in long-context tasks, highlighting their limited robustness. In contrast, our Fin-o1 models, trained on the FinCoT dataset, consistently outperform their backbone models and surpass larger general-purpose models, demonstrating the importance of domain-specific reasoning supervision. Further analysis shows that reinforcement learning methods have varying effectiveness in this context. While PPO and DPO fail to yield stable improvements, GRPO achieves consistent gains by aligning optimization with task-specific reasoning objectives. These findings illustrate that financial reasoning poses unique challenges that cannot be addressed by scaling model size or applying general reasoning techniques alone, emphasizing the need for specialized data, models, and optimization strategies.
2 FinCoT Reasoning Training Data
In this section, we describe the construction of FinCoT, the first high-fidelity dataset of chain-of-thought reasoning paths designed to enhance and evaluate LLMs’ financial reasoning capabilities, as illustrated in Fig. 1.
2.1 Data Collection
To develop a robust chain-of-thought training dataset tailored for financial reasoning, we systematically curate data from diverse financial data sources. Distinguished from FinR1 finr1 and Dianjin zhu2025dianjin , which only include FinQA for reasoning tasks while the rest of their datasets, i.e, Finance-Instruct flowers2025finance , FinCorpus duxiaoman2023fincorpus , and others that focus primarily on financial knowledge enhancement [^2] Samples of these data can be found in Appendix B.1, our dataset explicitly provides structured chain-of-thought supervision to enable complex financial reasoning, each contributing distinct reasoning skills [^3] See Appendix B for dataset details.:
  • FinQA chen2022finqadatasetnumericalreasoning and ConvFinQA chen2022convfinqa : Focused on numerical reasoning over short financial texts and single well-structured tables, supporting the model in basic arithmetic and aggregation tasks.
  • TATQA zhu2021tatqaquestionansweringbenchmark : Provides semi-structured financial reports with single-table QA tasks, enhancing the model’s ability to reason over less curated, real-world financial tables.
  • DocMath-Eval zhao2024docmath : Adds mathematical reasoning over structured tables, introducing complex quantitative reasoning patterns.
  • Econ-Logic quan2024econlogicqa : Includes financial decision-making and logical inference tasks, requiring models to apply domain-specific logic beyond numerical calculation.
  • BizBench-QA krumdick2024bizbench : Introduces multi-table reasoning tasks, training the model to integrate and synthesize information across multiple data sources.
  • DocFinQA reddy2024docfinqa : Comprises long-form financial documents with QA tasks, enabling the model to handle extended context and multi-step reasoning over lengthy narratives.
Figure 1: Overall framework of development of FinCoT, Fin-o1 and FinReason.
2.2 Iterative Chain-of-Thought Generation and Verification with Domain-Guided Supervision
Building our raw data collection, we further curate the first dataset containing 9,186 chain-of-thought annotated QA pairs, specifically designed to support SFT and RL for financial reasoning models, via a three-stage CoT generation framework by integrating domain-aware supervision, iterative refinement, and difficulty-aware filtering. This framework addresses the limitations of existing chain-of-thought methods, which prompt LLMs such as DeepSeek-R1 or Gemini team2024gemini ; ye2025limo ; muennighoff2025s1 to generate reasoning paths in a single round finr1 ; zhu2025dianjin , sufficient for general tasks but inadequate for financial reasoning which requires precise, multi-step, domain-informed calculations. For example, financial tasks like calculating profits require carefully following each calculation step with the right data and rules, where even small mistakes can completely invalidate the final answer.
In the first stage, we tackle the lack of structured guidance in existing methods by incorporating programmatic steps and expert-curated sub-questions from datasets such as BizBench and ConvFinQA, collected during our data preparation phase. While ConvFinQA offers high-quality sub-questions, it does not provide complete question-answer pairs that capture the full financial decision-making process. To close this gap, we use GPT-4o to synthesize these sub-questions into combined, holistic questions (see Fig. 2. This allows the model to see both the detailed steps and how they fit into a broader financial reasoning workflow, helping it generate more coherent and logically consistent reasoning paths that mirror domain expert expectations).
Figure 2: Workflow of curating the combined question and examples of program progress.
In the second stage, we address the brittleness of single-pass reasoning by introducing an iterative verification and refinement process, enabling the model to correct its reasoning progressively. Inspired by HuatuoGPT-o1 chen2024huatuogpt but adapted to the stricter demands of financial reasoning, we prompt GPT-4o to generate step-by-step reasoning paths, which are independently evaluated by a verifier (also GPT-4o). Instead of discarding incorrect outputs entirely, our framework forces the model to revisit earlier steps, refine its assumptions, and adjust its reasoning flow based on verifier feedback (see Fig. 3 in appendix C.). This allows the model to detect and fix errors that might otherwise propagate unchecked, ensuring greater accuracy and robustness, especially in complex financial tasks where early missteps can have outsized impacts.
In the third stage, we introduce a difficulty-aware filtering step to ensure that the final dataset targets genuinely challenging reasoning scenarios (as shown in Fig. 1). Unlike prior methods chen2024huatuogpt ; finr1 that focus only on verifying correctness, we also consider question difficulty, motivated by recent studies showing that easy QA pairs do little to advance reasoning ability ye2025limo ; muennighoff2025s1 . We use Llama3.1-8B-Instruct to attempt each question, then apply a verifier (see Section 4.3) to check the answers. QA pairs that are correctly solved by the filtering LLM are excluded, which is deemed less challenging.
3 Fin-o1
In this section, we propose Fin-o1 (including Fin-o1-8B and Fin-o1-14B) based on FinCoT, the first open source LLMs targeted on financial reasoning tasks. Qwen3 are selected as backbone models, given their strong performance among small and mid-sized LLMs (shown in Table 3). Following previous studies chen2024huatuogpt ; dubey2024llama , a two-stage training approach, supervised fine-tuning and RL, is adopted to develop the model. FinCoT is divided into two parts FinCoT_SFT (7686) and FinCoT_RL (1,500) to support the training. Based on FinCoT, we further conduct the first evaluation of different RL methods to assess their impact on financial reasoning tasks.
3.1 Supervised fine-tuning
In the first stage, we perform supervised fine-tuning, with the FinCoT_SFT, aiming to guide the model to systematically structure its reasoning process before providing the final answer. As shown in the right part of Fig. 1, each training instance is composed of a prompt, contextual information extracted from textual and tabular data, and a question that requires multi-step reasoning. The output target consists of a detailed reasoning path, followed by a summarized answer, emphasizing the importance of step-by-step logical analysis before deriving the final response. By structuring the output in this manner, the objective of SFT is to cultivate a disciplined reasoning framework, encouraging the model to first articulate a coherent analytical process, thereby enhancing interpretability and accuracy in complex financial reasoning scenarios. After the supervised fine-tuning, we get Fin-o1-SFT.
3.2 Reinforcement learning
In the second stage, we conduct a comprehensive evaluation of various RL methods to assess their impact on financial reasoning tasks. To effectively evaluate these methods, we specially design targeted reward functions and curate high-quality positive and negative samples tailored to each method. This comparative analysis aims to provide insights into the strengths and limitations of each RL method in enhancing reasoning capabilities for complex financial tasks.
Proximal policy optimization.
We start from PPO schulman2017proximal . It optimizes the policy by maximizing the expected reward while constraining the extent of policy updates, thus preventing excessive divergence from the reference policy. Each input $x$ is processed by the Fin-o1-SFT to generate a response with a reasoning path $y^{*}$. The reward $r$ is defined based on the verifier feedback when comparing if the $y^{*}$ is consistent with the gold answer $y$. To effectively evaluate the generated answer with the gold, GPT-4o is used to verify the correctness of the results.
$r(x,y,y^{*})=\begin{cases}1&\text{if }\text{verifier}(y,y^{*})=\text{True}\\ 0&\text{otherwise}\end{cases}$ (1)
Additionally, the reward is adjusted by incorporating the Kullback-Leibler (KL) divergence penalty between the updated policy $\pi_{\theta}$ and the reference policy $\pi_{ref}$ luong2024reft , aiming to prevent excessive deviation from the reference policy while ensuring stable learning.:
$r^{\prime}(x,y,y^{*})=r(x,y,y^{*})-\beta\cdot KL(\pi_{\theta}\parallel\pi_{ref})$ (2)
Direct preference optimization.
DPO rafailov2023direct refines the model’s decision-making by systematically contrasting positive and negative samples through pairwise comparisons. In our framework, we extend the conventional DPO setup by implementing a more strategically designed sampling mechanism tailored to the nuances of complex financial reasoning paths.
The iterative CoT reasoning path generation process is leveraged in this step to effectively curate high-quality positive-negative samples. As detailed in 2.2, this process involves employing GPT-4o to generate reasoning paths that undergo iterative verification and correction. Rather than employing low-quality negative samples that diverge significantly from the positive samples, we implement a targeted selection strategy that designates the last failed reasoning path $R_{n-1}$ as the negative sample. This path is specifically chosen as it closely mirrors the input query in logical structure while containing identifiable reasoning flaws, making it a more informative negative sample. By focusing on plausibility and logical coherence, our sampling strategy mitigates the risk of selecting trivial or overly simplistic negative samples, thereby reinforcing the model’s capacity to differentiate between nuanced reasoning paths and effectively learn from more challenging and contextually relevant negative samples.
Generalized proximal policy optimization.
Compared with PPO that optimizes the model by maximizing expected rewards through constrained policy updates, GRPO tang2024generalized extends PPO by incorporating additional objectives or multi-task learning capabilities, making it suited for financial reasoning tasks where accuracy and reasoning completeness, long-context reasoning abilities are crucial for effectively handling complex financial scenarios.
Therefore, to comprehensively improve LLMs’ reasoning abilities in financial reasoning tasks, we propose a novel multi-faceted reward structure in GRPO that systematically evaluates four critical dimensions. Compared with prior work that predominantly targets accuracy and format chen2024huatuogpt ; finr1 ; zhu2025dianjin , our framework introduces two additional components: reasoning logic and length rewards. By integrating these dimensions, we aim to mitigate shortcut learning, reinforce logical consistency, and incentivize effective reasoning over long contexts—a crucial capability given the inherently lengthy and complex nature of financial documents. Formally, the reward function for GRPO is designed as follows:
$r(x,y,y^{*})=\alpha_{1}\cdot r_{acc}+\alpha_{2}\cdot r_{logic}+\alpha_{3}\cdot r_{format}+\alpha_{4}\cdot r_{length}\cdot r_{acc}$ (3)
where:
  • $r_{acc}$: Accuracy reward, based on the verifier’s assessment of the final response correctness.
  • $r_{logic}$: Logical consistency reward, evaluating the correctness of the reasoning path.
  • $r_{format}$: Assessing if the generated response follows the designed format.
  • $r_{length}$: To encourage the model’s ability to reason over lengthy contexts—a critical requirement in financial analysis given the extensive nature of many financial documents—we provide an additional reward when the answer is correct and the context length exceeds 8,192 tokens. If the accuracy reward is zero, the length reward is also set to zero.
To effectively evaluate the generated response, we also leverage LLMs (GPT-4o) as judge. Following the reward score design in finr1 ; chen2024huatuogpt , we first assesses the accuracy of the generated response, signing a reward of 1 if the answer is deemed correct. Similarly, if the logical reasoning is validated as correct, a reward of 1 is also given. For format and length evaluation, we employ rule-based methods to assign corresponding rewards, the format reward is set as 0.1 and the length reward is set as 1.
4 FinReason Evaluation Benchmark
In this section, we present FinReason, the first comprehensive benchmark listed in Table 2 designed to systematically evaluate the financial reasoning capabilities of LLMs.
4.1 Tasks
Financial reasoning tasks present distinct challenges beyond general reasoning, as they require not only understanding the structure and semantic content of financial documents and tables but also accurately interpreting and executing complex equations. To this end, we introduce a unified benchmark that systematically integrates reasoning over (1) financial texts with tables, (2) long-context reasoning, (3) multi-table analysis, and (4) interpretation and execution of financial equations, establishing a comprehensive evaluation framework to assess and elevate LLM capabilities in financial reasoning tasks. Four datasets are selected to evaluate LLMs across these key aspects of financial reasoning:
(1) FinQA chen2021finqa serves as the foundational benchmark for quantitative reasoning over structured tables and basic financial texts. It assesses the model’s ability to extract and synthesize financial information related to core financial concepts such as revenue, profit margins, and financial ratios.
(2) Simplong zhao2024docmath targets long-context reasoning within single-table scenarios. It assesses the model’s capacity to perform numerical reasoning based on structured data while maintaining context throughout extended financial texts.
(3) Complong zhao2024docmath extends Simplong by introducing multi-table scenarios, challenging the model to integrate information across multiple tables. This dataset tests the model’s ability to handle more complex financial reporting contexts, where financial data is fragmented and requires cross-table reasoning to draw accurate conclusions.
(4) XBRL-Math wang2025finnlp emphasizes equation-based reasoning, requiring the model to accurately interpret and execute financial formulas derived from structured data in XBRL filings. This dataset assesses the model’s ability to associate financial terms with corresponding numerical data and perform structured calculations.
Table 2: Overview of the datasets evaluated in the study.
Dataset Size Data Types Data Source Avg. Token Metric License
FinQA chen2021finqa 1,100 Tables and Texts Financial reports 1,128 Acc MIT License
DM-Simplong zhao2024docmath 100 Tables and Texts Financial reports 4,330 Acc MIT License
DM-Complong zhao2024docmath 300 Tables and Texts Financial reports 39,983 Acc MIT License
XBRL-Math wang2025finnlp 90 Texts and Equations XBRL 397 Acc CDLA-Permissive 2.0
4.2 Evaluated Models
To investigate how existing LLMs, especially reasoning-enhanced LLMs, perform in financial tasks, we select a diverse set of 29 LLMs from the GPT, LLaMA, DeepSeek, and Qwen families, along with other reasoning models including both general domain and financial specific models. Our selection covers models of various scales, from smaller models (8B) to large-scale models (70B), and even ultra-large models like DeepSeek-R1 (671B). We also include advanced closed-source models such as GPT-4o and GPT-o1 as strong baselines. More details of the models can be found in Appendix E.
4.3 Evaluation Settings
In each dataset, LLMs are presented with a context containing financial reports, tables, equations, and symbols, along with a question. Following prior work wei2022chain , we use a generic prompt (see appendix G for more details) to instruct models to answer the question.
While answers to financial tasks are often expressed in mathematical forms, variations in representation, such as differences in percentages, rounding conventions, or numeric formatting, can lead to discrepancies in surface-level comparisons. To address this, we opt out exact match metrics and instead adopt an LLM-as-judge evaluation paradigm. Specifically, we follow the answer extraction methodology proposed by chen2023theoremqa in TheoremQA, extracting the numerical results from model outputs. The extracted answers are then compared against the ground truth to assess correctness, allowing for flexible yet rigorous evaluation of generated responses. Please see appendix D.2 for more details of the evaluation setting.
5 Results
Next, we present the results of these LLMs on FinReason and discuss our main findings through the results for the research questions listed in section  1.
Table 3: Overall performance of different LLMs on four financial reasoning datasets.
Models FinQA DM-Simplong XBRL-Math DM-Complong Average
GPT-4o 72.49 60.00 72.22 39.33 61.01
GPT-o1-preview 49.07 56.00 74.44 36.67 54.05
GPT-o3-mini 60.87 59.00 76.67 35.00 57.89
DeepSeek-V3 73.20 53.00 76.67 42.33 61.30
DeepSeek-R1 65.13 53.00 86.67 38.67 60.87
GPT-4.5 68.94 59.00 74.44 39.33 60.43
Llama-4-Scout 70.45 52.00 88.89 0.67 53.00
Llama-3-70B-Instruct 58.92 41.00 56.67 13.67 42.57
Llama-3.1-70B-Instruct 63.18 48.00 63.33 34.33 52.21
Llama-3.3-70B-Instruct 68.15 54.00 70.00 32.00 56.04
Qwen2.5-72B-Instruct 73.38 59.00 67.78 14.67 53.71
Qwen2.5-Math-72B-Instruct 69.74 42.00 83.33 5.00 50.02
DeepSeek-R1-Distill-Llama-70B 66.73 53.00 86.67 30.67 59.27
Qwen2.5-32B-Instruct 73.11 56.00 65.56 30.00 56.17
Qwen3-32B 64.15 51.00 85.56 26.00 56.68
QwQ-32B 61.22 46.00 84.44 20.00 52.92
DeepSeek-R1-Distill-Qwen-32B 65.48 55.00 84.44 24.67 57.40
Limo 63.44 45.00 61.11 15.33 46.22
S1-32B 66.81 53.00 84.44 24.00 57.06
Qwen2.5-14B-Instruct 67.44 59.00 57.78 26.67 52.72
Qwen3-14B 64.33 49.00 86.67 24.00 56.00
DeepSeek-R1-Distill-Qwen-14B 63.27 44.00 84.44 21.00 53.18
DeepSeek-R1-Distill-Llama-8B 45.96 33.00 81.11 15.67 43.94
Llama-3-8B-Instruct 41.97 29.00 48.89 6.00 31.47
Llama-3.1-8B-Instruct 54.13 34.00 62.22 14.30 41.16
Qwen2.5-7B-Instruct 55.37 41.00 42.22 17.67 39.07
Qwen3-8B 62.11 46.00 83.33 17.67 52.28
FinR1-7B 58.74 37.00 30.00 13.67 34.85
Dianjin-R1-7B 60.20 35.00 83.33 14.67 48.30
Fino1-8B 73.03 56.00 84.44 26.33 59.95
Fino1-14B 74.18 55.00 87.78 27.33 61.07
Table 3 presents the performance of evaluated LLMs across four financial reasoning benchmarks: FinQA, DM-Simplong, XBRL-Math, and DM-Complong. Our Fin-o1-14B achieves a score of 61.07, ranking second only to Deepseek-V3 despite having substantially fewer parameters. Notably, Fin-o1-14B surpasses larger-scale models like LLaMA-3.3-70B-Instruct, Qwen2.5-72B-Instruct, even GPT-o1-preview, GPT-o3-mini, and GPT-4.5, underscoring its superior financial reasoning capabilities. Fin-o1-8B also performs competitively, attaining an average score of 58.58, exceeding most models with 70B parameters or fewer. Both Fin-o1-8B and Fin-o1-14B excel in FinQA, DM-Simplong, and XBRL-Math, demonstrating the impact of targeted domain adaptation in financial reasoning tasks. These gains also confirm the effectiveness of our three-stage FinCoT construction and the two-phase SFT + RL training strategy.
General reasoning LLMs fail in financial reasoning.
While general reasoning models exhibit significant improvements in general mathematical reasoning tasks, they struggle to effectively transfer these gains to financial reasoning tasks, which demand specialized knowledge and reasoning over financial documents and tables. For instance, the performance of QwQ-32B and Limo declined notably compared to Qwen2.5-32B-Instruct, dropping from 56.17% to 52.92% and 46.22%, respectively. Similarly, although S1-32B demonstrated overall performance gains, its effectiveness in FinQA, Simplong and Complong decreased. These results highlight the persistent gap between general and financial reasoning due to the complexities of interpreting financial tables and structures.
Incomprehensive reasoning data limits consistent improvements. While our model demonstrates consistent improvements over the backbone model Qwen3 across all datasets, other financial models do not exhibit such stable improvements. For instance, while Fin-R1 and Dianjin-R1-7B achieved higher scores in FinQA, their performance in long-context reasoning tasks dropped significantly compared with their backbone model Qwen-2.5-7B-Instruct. On Simplong, the backbone model scored 41%, but Fin-R1 and Dianjin-R1-7B fell to 37% and 35%, respectively. A similar decline is observed on Complong, indicating the challenges in maintaining long context reasoning performance. These declines are attributable to CoT datasets used in these works, which focused on shorter contexts and knowledge-based QA rather than reasoning over contexts, limiting the models’ exposure to reasoning over extended financial contexts.
Table 4: Ablation study of Fin-o1-8B with different SFT training data and RL methods.
Models FinQA DM-Simplong XBRL-Math DM-Complong Average
Backbone model
Qwen3-8B 62.11 46.00 83.33 17.67 52.28
SFT model
Fin-o1-8B (SFT with original QA pair) 66.55 50.00 74.44 16.67 51.92
Fin-o1-8B (SFT with all CoT data) 71.32 51.00 81.11 21.67 56.28
Fin-o1-8B (SFT with filtered CoT) 71.78 57.00 82.22 23.33 58.58
SFT + RL model
Fin-o1-8B (SFT + PPO) 72.05 54.00 83.33 23.00 58.10
Fin-o1-8B (SFT + DPO) 69.74 51.00 72.22 25.00 54.49
Fin-o1-8B (SFT + GRPO) 73.03 56.00 84.44 26.33 59.95
Table 4 presents additional results evaluating the effectiveness of various RL methods in financial reasoning tasks and further verifying the effectiveness of FinCoT. Based on the results, we find that:
GRPO Outperforms Other RL Methods in Financial Reasoning.
GRPO achieves the most substantial gains, driven by its multi-faceted reward function that emphasizes correct reasoning and long-context comprehension. Notably, in DM-Complong, where other models like Fin-R1 and S1-32B falter, the SFT+GRPO model reaches 26.33% with just 8B parameters—a 9% improvement over its backbone model. In contrast, PPO shows gains in FinQA and XBRL-Math but declines in Simplong and Complong, indicating weaker long-context reasoning. For DPO, while the performance in DM-Complong improves, performance in simpler tasks declines. This may stem from the design of positive-negative sample pairs, where the negative path is curated from the last-mistake path and the positive path is generated by allowing the model to take one additional reasoning step. This approach may inadvertently encourage excessive reasoning, benefiting complex tasks but potentially causing overthinking in simpler tasks.
We further compare the impact of fine-tuning the backbone model using original QA pairs without CoT reasoning and with unfiltered CoT data. Training without CoT data slightly improves basic financial reasoning tasks like FinQA, but leads to performance drops in XBRL-Math and Complong. This decline is particularly pronounced in XBRL-Math, where the lack of step-by-step reasoning impedes the model’s ability to handle financial equations, hindering its generalization to equation-heavy tasks. In contrast, using the unfiltered FinCoT dataset demonstrates the effectiveness of exposing the model to more complex reasoning paths, underscoring the importance of reasoning-focused training for financial tasks.
6 Conclusion
In this work, we address the critical yet underexplored challenge of financial reasoning for LLMs. We present three key contributions to bridge existing gaps: (1) the development of FinCoT, the first publicly available CoT dataset specifically curated for financial reasoning tasks; (2) the introduction of Fin-o1, the first open-source financial reasoning model trained through a two-phase framework combining supervised fine-tuning and RL methods; and (3) the establishment of FinReason, a comprehensive benchmark designed to systematically evaluate financial reasoning across essential components such as multi-table analysis, long-context reasoning, and equation-based tasks.
Our extensive evaluations demonstrate that general reasoning models struggle to adapt to financial contexts, revealing significant overall performance drops. In contrast, our Fin-o1 models, trained on FinCoT with GRPO, achieve consistent gains across diverse financial reasoning tasks, highlighting the critical role of domain-specific CoT supervision and structured RL optimization. These findings underscore the need for targeted datasets, specialized models, and advanced training strategies to effectively address the complexities of financial reasoning, setting a new standard for future research in this domain.
Appendix A Related Work
Our work builds upon three main research directions: the development of general-purpose and reasoning-enhanced language models, applications of LLMs in finance, and studies on reasoning capabilities in financial tasks.
A.1 General-Purpose and Reasoning LLMs
Recent years have witnessed significant advancements in large language models which demonstrated remarkable capabilities across various tasks, including natural language understanding, generation, and complex reasoning. Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting [35], self-reflection [44], and reinforcement learning from human feedback [23]. These models have shown superior performance in mathematical reasoning [16], symbolic manipulation [7], and logical inference tasks [2].
A.2 LLMs in Financial Applications
The application of LLMs in finance has gained significant attention, with models being adapted for tasks such as financial text analysis [42], market sentiment prediction [41], and automated trading strategies [13]. Recent work has focused on developing finance-specific models like BloombergGPT [36], FinGPT [41] and PIXIU [38], OpenFinLLMs [39], Plutus[24], which are pretrained on financial corpora to better understand domain-specific terminology and concepts. These models have demonstrated effectiveness in tasks such as financial sentiment analysis, earnings call analysis, and regulatory compliance checking.
A.3 Financial Reasoning
Financial reasoning presents unique challenges that combine numerical computation, domain knowledge, and logical inference. Recent studies have explored LLMs’ capabilities in financial reasoning tasks, including numerical reasoning over financial documents [4], mathematical problem-solving in financial contexts [29], and structured data interpretation [18]. These works highlight the importance of both domain expertise and reasoning abilities in financial applications. Our work extends these research directions by systematically evaluating how reasoning enhancements in state-of-the-art LLMs translate to financial domain tasks, providing insights into the effectiveness of different reasoning strategies and identifying areas for improvement in financial reasoning capabilities. Additionally, we focus on benchmarking these models across diverse financial datasets, emphasizing multi-table reasoning, long-context comprehension, and quantitative analysis. This comprehensive evaluation not only sheds light on current model limitations but also informs future research directions for developing more robust financial reasoning models.
Appendix B Details of Data Source
B.1 Data source examples for other works
This section present some examples of the data from other models. In convince of reviewing, we have translated all the original context into English if they are not English.
Finance-Instruct Explain tradeoffs between fiscal and monetary policy as tools in a nations economic toolkit. Provide examples of past instances when each were utilized, the economic conditions that led to them being deployed, their intended effects, and an evaluation of their relative efficacy and consequences.
CFLUE Assume you are a financial industry expert. Please answer the following question. Note: The question is a single-choice question. Only return the most appropriate option. If multiple options are suitable, only return the most accurate one. Among the following options, which financial institution is not permitted to participate in the interbank lending market? A. Clearing Center B. Insurance Company C. Securities Company D. Fund Management Company
CFLUE Is the sentiment of this text positive, negative, or neutral? Feeling exhausted from averaging down.
Finance-IQ When considering the pricing of life insurance products, several key factors come into play, including market forces, business operating costs, and government regulations. Among these factors, which one primarily determines the minimum limit of life insurance product prices? A. Market forces and business operating costs B. Business operating costs C. Government regulations D. Market forces
B.2 Data source for FinCoT
We provide here detailed descriptions of the datasets used to construct our training and evaluation corpus for financial reasoning enhancement.
  • FinQA [5]: FinQA consists of over 5,000 open-ended question-answer pairs derived from financial reports and tables. It integrates both textual and tabular data, covering numerical reasoning tasks such as arithmetic operations, comparison, and multi-step calculations, making it well-suited for training reasoning-enhanced models.
  • DocFinQA [27]: An extension of FinQA designed for long-context reasoning. It pairs FinQA-style questions with full-length financial filings, increasing the average context length from under 700 words to approximately 123,000 words. This dataset better simulates real-world document scale and complexity, enabling robust evaluation of retrieval and reasoning over long financial documents.
  • TAT-QA [47]: A large-scale financial QA dataset created from real financial reports. It combines both textual and tabular data and emphasizes diverse numerical reasoning tasks, including arithmetic operations, counting, sorting, and comparison. TAT-QA reflects real-world challenges in financial document interpretation.
  • ConvFinQA [6]: A financial question answering dataset designed to handle multi-turn conversations involving financial tables and text. ConvFinQA builds upon FinQA by incorporating follow-up questions and context retention across multiple turns, simulating more realistic financial analysis scenarios. It provides expert-annotated sub-questions that decompose complex financial questions into granular steps, facilitating the generation of structured reasoning paths and promoting systematic financial reasoning.
  • DocMath-Eval [45]: This benchmark suite evaluates mathematical reasoning in domain-specific contexts. We utilize two subsets:
    • DocMath-CompShort[45]: Constructed from quarterly and annual reports, it involves reasoning across long documents and multiple tables, presenting a challenging multi-step numerical reasoning setting.
    • DocMath-SimpShort[45]: Built from FinQA and TAT-QA, this subset focuses on localized numerical reasoning tasks over short documents containing a single table.
  • Econ_Logic [25]: It consists of financial logic reasoning tasks that help the model to further understand and have more insights of financial terms and knowledge.
  • BizBench-QA [14]: Contains financial QA pairs across multiple task types, including SEC-NUM, CodeFinQA, CodeTAT-QA, and FinCode. These tasks span code generation, numerical span identification, and financial mathematical reasoning from various perspectives, enriching the training corpus with broader financial task coverage.
FinQA Please answer the given financial question based on the context. Context: measurement point december 31 the priceline group nasdaq composite index s&p 500 rdg internet composite . |measurement pointdecember 31|the priceline group inc .|nasdaqcomposite index|s&p 500index|rdg internetcomposite| |2010|100.00|100.00|100.00|100.00| |2011|117.06|100.53|102.11|102.11| |2012|155.27|116.92|118.45|122.23| |2013|290.93|166.19|156.82|199.42| |2014|285.37|188.78|178.29|195.42| |2015|319.10|199.95|180.75|267.25|. Question: what was the percent of the growth of the the priceline group inc . from 2014 to 2015 Answer:
Appendix C Example of CoT data curation
Here is an example of applying the guideline and iterative refinement process to generate a CoT reasoning path based on the given context and question.
Figure 3: Example of iterative refinement of the CoT process with a guidance.
BizBench From the results reflected in the following pro forma information because of normal production declines, changes in commodity prices, future acquisitions and divestitures, future development and exploration activities and other factors. | | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | | | (In millions, except per share amounts) | | Six Months Ended June 30, 2021 | | Pro forma revenue | | \$ | 1,802 | | | | | Pro forma net income | | 71 | | | | | | | | | | | | | | | | | | | | | | | | | | | | | Pro forma basic earnings per share | | \$ | 0.09 | | | | | Pro forma diluted earnings per share | | \$ | 0.09 | | | | 3. Properties and Equipment, NetProperties and equipment, net are comprised of the following: | | | | | | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | (In millions) | | June 30,2022 | | December 31,2021 | | Proved oil and gas properties | | \$ | 16,102 | | | \$ | 15,340 | | | Unproved oil and gas properties | | 5,292 | | | 5,316 | | | Gathering and pipeline systems | | 423 | | | 395 | | | Land, buildings and other equipment | | 144 | | | 140 | | | Finance lease right-of-use asset | | 24 | | | 20 | | | | | 21,985 | | | 21,211 | | | Accumulated depreciation, depletion and amortization | | (4,578) | | | (3,836) | | | | | \$ | 17,407 | | | \$ | 17,375 | | As of June 30, 2022, the Company had no projects with exploratory well costs capitalized for more than one year after drilling. | | | | | | | | | | | | | | | | | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | --- | | | | | | | | | | | | | | | | | | (In millions) | | June 30,2022 | | December 31,2021 | | 6.51% weighted-average private placement senior notes (1) | | \$ | 37 | | | \$ | 37 | | | 5.58% weighted-average private placement senior notes (2) | | 87 | | | 87 | | | 3.65% weighted-average private placement senior notes | | 825 | | | 825 | | | 4.375% senior notes due June 1, 2024 | | 750 | | | 750 | | | 3.90% senior notes due May 15, 2027 | | 750 | | | 750 | | | 4.375% senior notes due March 15, 2029 | | 500 | | | 500 | | | Revolving credit facility | | - | | | - | | | Net premium (discount) | | 164 | | | 185 | | | Unamortized debt issuance costs | | (8) | | | (9) | | | | | \$ | 3,105 | | | \$ | 3,125 | | (1) Includes \$37 million of current portion of long-term debt at June 30, 2022.(2) Includes \$87 million of current portion of long-term debt at June 30, 2022.At June 30, 2022, the Company was in compliance with all financial and other covenants for both its revolving credit facility and senior notes. Question: What is the pro forma net income?
Appendix D Details of experiments
D.1 Model training
We list the details of the training process here. The SFT process is configured with a learning rate of 5e-6 and a weight decay of 0.1. Training is conducted for 3 epoch with a batch size of 2 per GPU and a gradient accumulation of 2 steps, effectively resulting in an effective batch size of 2. The gradient checkpointing is enabled to optimize memory usage. For RL, the training process is configured with a learning rate specified set to 5e-6. The gradient accumulation is set to 2 steps, and a gradient accumulation of 2 steps, effectively resulting in an effective batch size based on the configured gradient_accumulation_steps. The training is conducted using 3 GPUs, each handling a separate training process. And a total of 500 steps is used to train the model.
D.2 Model evaluation
For large models (e.g., DeepSeek-V3, DeepSeek-R1, GPT series, Llama4), we run evaluations via APIs from TogetherAI [^4] https://www.together.ai and OpenAI [^5] https://platform.openai.com, using versions GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06. All open-sourced models up to 72B parameters were conducted with vLLM [15] framework on 8 A100 GPUs (80GB each), with a max generation length of 1,024 tokens and temperature setting of 0. Given the extreme length of some evaluated dataset (e.g., Dm-Complong with a average length of 40K tokens), we use a zero-shot setting, aim to evaluating LLMs’ capabilities for financial reasoning without any few-shot demonstrations.
Appendix E Details of Evaluated Models
To comprehensively evaluate the performance of LLMs in financial tasks, particularly those with reasoning enhancements, we select a diverse set of 29 LLMs. This selection includes both general-purpose, and finance-specific models, ranging from smaller models (8B) to large-scale models (70B), and extending to ultra-large models such as DeepSeek-R1 (671B), GPT-4o, and GPT-o1. The evaluated models are summarized as follows: GPT-4o [11]: OpenAI’s flagship model with multimodal capabilities and improved real-time reasoning. GPT-o1 [12]: A reasoning-optimized variant in the GPT series with strong step-by-step problem-solving ability. GPT-o3-mini: A compact version of GPT-o3, optimized for efficiency with robust language understanding. DeepSeek-V3 [17]: A 671B Mixture-of-Experts (MoE) model with 37B active parameters per token, designed for efficient reasoning. DeepSeek-R1 [10]: A first-generation reasoning model with competitive performance on math, code, and reasoning tasks. Qwen series [40]: A strong multilingual model with advanced instruction tuning. LLaMA-series [8]: Widely adopted open-source models (8B–70B) as well as the latest MOE models with strong instruction-tuned variants used in our experiments. DeepSeek-R1-Distill [10]: Distilled models from DeepSeek-R1, fine-tuned using LLaMA and Qwen backbones to balance performance and efficiency. Financial-Specific Reasoning Models: This category includes models such as Fin-R1-7B [19] and Dianjin-R1-7B [48], both of which leverage financial reasoning datasets and employ fine-tuning and reinforcement learning techniques to optimize LLM performance in financial reasoning tasks. Other reasoning models: They include S1-32B [22], Limo [43]. These two models are designed for general reasoning tasks, with limited but high quality COT data.
More details of the evaluated models are presented in Table 5. It also highlights the diverse training strategies employed across the evaluated models. Reinforcement learning has been extensively utilized to enhance reasoning capabilities, with common methods including PPO, DPO, and GRPO. Among these, GRPO is the most widely adopted, as seen in models such as DeepSeek-R1, Qwen2.5-72B-Instruct, and FinR1-7B. This widespread use of GRPO underscores its effectiveness in refining reasoning abilities, particularly in complex, multi-step tasks involving financial documents.
Table 5: Summary of evaluated LLMs, including their parameter sizes, reasoning capabilities, input limits, source availability, and reasoning enhancement strategies. MCTS refers to Monte Carlo Tree Search. CoT refers to chain-of-thought reasoning. TIR means tool-integrated reasoning.
Model Name Parameters Training Window Size Close/Open Source Reasoning Training Data
GPT-4o - - 128k Closed -
GPT-o1 - RL with long CoT 128k Closed Public and proprietary data (Human-annotated CoT, MCTS-assisted Synthetic data)
GPT-o3-mini - - 128k Closed Public and proprietary data
DeepSeek-V3 671B GRPO 128k Open -
DeepSeek-R1 671B GRPO 128k Open Cold-start data generation, post-processing data
GPT-4.5 - - 128k Closed -
Llama4-Scout 17B*16E SFT & DPO 128k Open -
Qwen2.5-72B-Instruct 72B DPO & GRPO 128k Open -
Qwen2.5-72B-Instruct-Math 72B DPO & GRPO 128k Open Synthetic data from Qwen, high-quality mathematical data, CoT, TIR
DeepSeek-R1-Distill-Llama-70B 70B GRPO 128k Open Distilled from R1
Llama3-70B-Instruct 70B SFT & DPO 8k Open -
Llama3.1-70B-Instruct 70B SFT & DPO 128k Open -
Llama3.3-70B-Instruct 70B SFT & DPO 128k Open -
Qwen2.5-32B-Instruct 32B DPO & GRPO 128k Open -
Qwen3-32B 32B RL with long CoT 128k Open -
DeepSeek-R1-Distill-Qwen-32B 32B GRPO 128k Open Distilled from R1
Limo 32B SFT 128k Open Public data with Qwen filteration, with CoT generated from Deepseek
S1-32B 32B SFT 128k Open Public data and original data, filtered by quality, with CoT generated from Gemini
Qwen/QwQ-32B 32B PPO 128k Open -
Qwen2.5-14B-Instruct 14B DPO & GRPO 128k Open -
Qwen3-14B 14B RL with long CoT 128k Open -
DeepSeek-R1-Distill-Qwen-14B 14B GRPO 128k Open Distilled from R1
DeepSeek-R1-Distill-Llama-8B 8B GRPO 128k Open Distilled from R1
Llama3-8B-Instruct 8B SFT & DPO 8k Open -
Llama3.1-8B-Instruct 8B SFT & DPO 128k Open -
Qwen2.5-7B-Instruct 7B DPO & GRPO 128k Open -
Qwen3-8B 8B - 128k Open -
FinR1-7B 7B SFT & GRPO 128k Open 60k CoT data for financial tasks
Dianjin-R1-7B 7B SFT & GRPO 128k Open 31k CoT data for financial tasks
Figure 4: Error case 1.
Figure 5: Error case 2.
Appendix F Error Analysis
To further investigate why reasoning-enhanced models like DeepSeek-R1 and GPT-o1 underperform in financial tasks compared to general-purpose models such as GPT-4o and DeepSeek-V3, we conduct an error analysis using the FinQA dataset since these reasoning-enhanced models exhibit a significant performance drop on FinQA.
We focus on cases where DeepSeek-R1 produces incorrect answers while DeepSeek-V3 generates correct ones. Two common error patterns emerge. First, DeepSeek-R1 tends to over-reason and does not strictly adhere to instructions, as illustrated in Figure 4. In this case, the question requires a sum, but the model provides excessive details by breaking down the calculation instead of directly summing the values as instructed.
Second, compared to DeepSeek-V3, DeepSeek-R1 appears to lack financial sensitivity. For instance, in Figure 5, it fails to recognize critical financial nuances, leading to an incorrect response. Specifically, DeepSeek-R1 misinterprets the concept of "average share price" in the UK Employee Share Purchase Plan by incorrectly identifying it as the employee purchase price (\$4,078 per share) rather than the actual compensation expense per share (\$719 per share). This mistake stems from an inability to correctly differentiate between compensation expense per share and the full fair market value (FMV) calculation, resulting in an overestimated figure. In contrast, DeepSeek-V3 correctly understands that the compensation expense per share represents the direct financial impact per issued share, allowing it to provide the accurate answer. This highlights a fundamental gap in DeepSeek-R1’s financial reasoning capabilities, particularly in recognizing the accounting conventions used to report share-based compensation.
Appendix G Prompt
Prompt Template for DocMath-CompLong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context. You need to first think through the problem step by step, documenting each necessary step. Then you are required to conclude your response with the final answer in your last sentence as Therefore, the answer is {final answer}’. The final answer should be a numeric value. ###Context <Full document here> ### Input <the question based on the document> Lets think step by step to answer the given question. ### Output
Prompt Template for DocMath-Simplong task You are a financial expert, you are supposed to answer the given question based on the provided financial document context. You need to first think through the problem step by step, documenting each necessary step. Then you are required to conclude your response with the final answer in your last sentence as Therefore, the answer is {final answer}’. The final answer should be a numeric value. ###Context <simple long context here> ### Input <the question based on the context> Lets think step by step to answer the given question. ### Output
Prompt Template for FinQA task Please answer the given financial question based on the context. Context: <context> Question: <the question based on the context> Answer:
Prompt Template for XBRL-Math task You are a financial expert tasked with carefully reading, analyzing, and answering the following eXtensible Business Reporting Language. Please follow the steps below: INPUT: Read the eXtensible Business Reporting Language (XBRL) question: <XBRL related question>, formula: <the formula related to this question>, and the explanation: <term explanation>. Provide only the final answer which is the numerical result of the calculation. For formulas like ROI, provide percentages. Never use the percent symbol in percentages. OUTPUT:
Limitations
Our study has several limitations that highlight areas for future improvement. First, our model scale is limited, as we only fine-tuned an 8B model (Fino1), while larger models (e.g., 70B) could potentially benefit more from reasoning enhancements. Second, our evaluation scope is restricted, covering only three financial reasoning tasks (FinQA, DM-Simplong, and XBRL-Math), which do not fully capture the breadth of financial NLP applications such as forecasting, financial sentiment analysis, and fraud detection. Third, our fine-tuning relies on a single dataset, FinQA, for reasoning path construction, limiting the model’s exposure to different financial reasoning patterns; incorporating additional datasets could improve generalization. Fourth, our reasoning path construction approach is simplified, as we generate paths using a single method (GPT-4o), whereas exploring multiple reasoning path generation strategies—such as ensemble approaches or human-annotated paths—could lead to more robust financial reasoning capabilities.
Future Direction
Developing effective reasoning models for financial tasks poses distinct challenges compared to general-domain reasoning. Our findings suggest three key directions for improvement: (1) Enhancing financial knowledge and terminology understanding through domain-specific pretraining using financial corpora, structured reports, and regulatory documents; (2) Improving multi-table reasoning and long-context comprehension by developing strategies that better infer logic across tables and optimize memory mechanisms for processing lengthy financial documents; and (3) Refining reasoning-enhanced strategies, as current approaches like GPT-o1’s chain-of-thought prompting and DeepSeek-R1’s self-reflection via reinforcement learning yield mixed results—highlighting the need for more tailored reasoning strategies to address both structured numerical and unstructured textual financial reasoning.
Ethics Statement
While our study demonstrates significant advancements in financial reasoning with LLMs, Fino1 remains a research prototype and is not yet suitable for real-world financial applications. Our model inherits the well-documented limitations of large language models, including hallucinations, sensitivity to input phrasing, and potential biases in financial data sources, which could affect reliability in critical financial contexts. Furthermore, Fino1 has not undergone rigorous testing for high-stakes financial decision-making, such as investment strategies, regulatory compliance, or risk assessment, where incorrect outputs could lead to severe consequences. Additionally, the model’s reliance on a single dataset (FinQA) and a single reasoning path construction method (GPT-4o) limits its robustness and adaptability to the complex and evolving nature of financial environments. As with all AI-driven financial models, careful human oversight remains essential, and future work must focus on enhancing factual consistency, mitigating biases, improving domain adaptation, and ensuring alignment with regulatory frameworks before such models can be considered for real-world financial deployment.
Potential impacts
The proposed work presents both positive and negative societal implications. The development of advanced financial reasoning models has the potential to enhance financial analysis by providing more accurate predictions and robust risk assessments, thereby supporting informed decision-making in areas such as investment strategies and regulatory compliance. However, these models may also pose risks if misused, such as reinforcing existing biases in financial datasets or facilitating manipulative financial practices. To mitigate such risks, we incorporate safeguards that enforce controlled access to the models, implement guidelines for responsible use, and maintain transparency in model outputs. Additionally, we emphasize the importance of adhering to ethical guidelines throughout the development and deployment process to minimize unintended societal consequences.